# Supervised Manifold Learning for Functional Data

### v1

@@ 5 Simulation Studies
We conduct simulation experiments to evaluate performance of the functional supervised manifold learning (FSML) procedure on finite samples. To construct a functional classifier, we combine the FSML with three classical classifiers: the $k$-NN classifier (FSMLk-NN), the SVM (FSMLSVM), and the linear discriminant analysis (FSMLLDA). For comparison, we consider several existing functional classifiers: the centroid classifier (CC, Delaigle and Hall,, 2012), the functional quadratic discriminant analysis (FQDA, Delaigle and Hall,, 2013; Galeano et al.,, 2015), the nonparametric Bayesian classifier (NB, Dai et al.,, 2017) based on the kernel density estimate, and the functional deep neural network classifier (FDNN, Wang et al.,, 2023). The code to implement the CC and FDNN are available on the authors’ personal website, while the FQDA and NB are reproduced using our code.
We generate data from a few functional manifold models as well as a Gaussian process model:
@@
    Torus: $X_{i}(t)=\{2+\cos(\theta_{i})\}\cos(\phi_{i})\sin(2\pi t)+\{2+\cos(\theta_{i}\}\sin(\phi_{i})\cos(2\pi t)+\sin(\theta_{i})\sin(4\pi t)$, where $(\theta_{i},\phi_{i}|Y_{i}=0)$ follows a uniform distribution on $0<\theta<\phi<2\pi$, and $(\theta_{i},\phi_{i}|Y_{i}=1)$ follows a uniform distribution on $0<\phi\leq\theta<2\pi$.
@@
    Gaussian: $X_{i}(t)=\xi_{i1}\log(t+2)+\xi_{i2}t+\xi_{i3}t^{3}$, where $(\xi_{i1},\xi_{i2},\xi_{i3})^{\top}|Y=k\overset{iid}{\sim}N(\mu_{k},\Sigma_{k})$ with $\mu_{0}=(-1,2,-3)^{\top},\mu_{1}=(-1/2,5/2,-5/2)^{\top},\Sigma_{0}^{1/2}=\operatorname{diag}(3/5,2/5,1/5)$ and $\Sigma_{1}^{1/2}=\operatorname{diag}(9/10,1/2,3/10)$.
Under all models above, the label variable $Y_{i}\overset{iid}{\sim}b(1,0.5)$ corresponding to a binary classification setting. The observed data are $(\widetilde{X}_{ij},Y_{i})$, where $\widetilde{X}_{ij}=X_{i}(t_{j})+\epsilon_{ij}$, for $i=1,\ldots,n$ and $j=1,\ldots,J$. Here, $10=t_{1}<\ldots<t_{J}=$ 1 are equidistant points on $[0,1]$, and $\epsilon_{ij}\overset{iid}{\sim}N(0,\widehat{V}_{X}/20)$, where $\widehat{V}_{X}$ is the sample variance of $X$ integrated on $[0,1]$.
Model (i), modified from model (iii) in Tan et al., (2024), is a time-warping model exhibiting significant phase variation, which is a classical example of a functional manifold (Srivastava et al.,, 2011; Chen and Müller,, 2012). Models (ii) and (iii) are two connected Swiss rolls (see the left panel of Figure 1 for an illustration) and a torus embedded into a functional space, respectively. These geometric objects are two-dimensional manifold examples commonly used in the literature on manifold learning (Tenenbaum et al.,, 2000; Meilă and Zhang,, 2024). Model (iv) is a Gaussian process model adopted from Model 1 in Wang et al., 2024b .
@@
Model $n$ FSML20-NN FSMLSVM FSMLLDA CC FQDA NB FDNN
@@
We generate training data following each of models (i) to (iv) above with two sampling settings $(n,J)=(200,50)$ and $(500,100)$, while an independent sample of size 500 is generated for testing. The ridged local linear estimator with the plug-in bandwidth as in Section 3.1 is applied to obtain smooth curves before applying any method. We set $k_{\rm{PCA}}=15$ and apply the nested ten-fold CV procedure described in Section 3.4 to select $\xi$ and $h_{\rm{reg}}$. For any method involving FPCA, we choose the number of PCs to explain at least 95% of the variance. We apply all the methods to each of the settings repeatedly for 200 times, and the misclassification errors are summarized in Table 1. We see that the CC performs the best under model (i), followed closely by the NB and FSML methods. The FDNN performs poorly under model (i). Under models (ii) to (iv), the FSML methods significantly outperform the benchmark methods. In particular, the FSML methods achieve near perfect classification under model (ii), where the CC and FQDA perform poorly. Also, it is worth noting that our methods still perform the best under the Gaussian process model (iv), although we assume nothing about the underlying data distribution. A possible reason is that our method is particularly efficient if the intrinsic dimension is low, while the CC, FQDA and NB are more powerful when the intrinsic dimension is high, as their asymptotic perfect classification results require the number of PCs tends to infinity. A similar observation has been noted in Wang et al., 2024b . Finally, we note that the differences among the three FSML methods are tiny, indicating that the classification performance of the FSML is insensitive to the chosen specific multivariate classifier.
@@
One of the applications of FSML is data visualization. To illustrate this, we generate two samples from models (ii) and (iii), respectively, and apply the FSML, as well as the FPCA for comparison. We show in Figure 3 the smoothed data by the ridged local linear estimator, the first two PCs by the FPCA, and 2-dimensional manifold learning outcomes by FSML with several $\xi$’s. We see that the FSML properly unfold the Swiss rolls and the outcomes can be linearly separated according to their classes (even with $\xi=0$) under model (ii). Similarly, the FSML outcomes from different classes are also separable with $\xi=4$ under model (iii). In contrast, the first two PCs do not linearly separate two classes under both of the models.
@@ 6 Real Data Applications
The first dataset concerns gene expression levels of a yeast cell based on the $\alpha$ factor experiment (Spellman et al.,, 1998, data available at the supplemental material). Several thousand genes were measured every seven minutes from 0 to 119 minutes, of which 612 genes were found to fluctuate periodically within the cell cycle. We focus on the subset that exhibits significant periodic features: the trajectories belonging to the $G_{1}$ group are coded with $Y=0$ and the trajectories belong to the $G_{2}$ and $M$ groups are coded with $Y=1$. We then obtain a sample of 473 trajectories, among which 223 (resp., 250) trajectories belong to $Y=0$ (resp., $Y=1$). Because the observations are quite noisy and the number of observations per gene is moderate, we use the local linear estimator with a manually chosen large bandwidth to recover smooth functional data.
@@
The raw data for these three datasets are presented in Figure 4. To evaluate the classification performance, we randomly partition each of the above datasets into a training sample (80% of the dataset) and a testing sample (20% of the dataset). Each of the functional classifiers is trained using the training sample and evaluated on the testing sample. The procedure is repeated 100 times, and the mean and standard deviation of the misclassification errors are reported in Table 2. To implement the FSML methods, the number $k_{\rm{PCA}}$ is set to 10, the intrinsic dimension $d$ is estimated by the minimal neighborhood method (Facco et al.,, 2017) summarized in Section B of the Supplementary Material, and $\xi$ and $h_{\rm{reg}}$ are chosen following the same criterion as in Section 5.
We see from Table 2 that, in the yeast gene dataset, the CC performs the best with a little advantage over the FSML methods, followed by the FQDA and NB, while the FDNN performs less well. Under the DTI dataset, the FSML methods clearly outperform the others. Finally, the FSML5-NN performs the best under the wine dataset, closely followed by the NB and the other two FSML methods. Overall, the FSML methods perform outstandingly under these three datasets.
Table 2: Mean (standard deviation) misclassification errors in percentage under all real datasets with the best result highlighted in boldface.
Dataset FSML5-NN FSMLSVM FSMLLDA CC FQDA NB FDNN
Yeast gene 6.0 (2.5) 5.8 (2.4) 5.7 (2.4) 5.4 (2.0) 7.3 (2.4) 7.6 (2.5) 15.2 (19.3)
DTI 22.1 (6.7) 22.6 (6.5) 22.5 (7.0) 26.0 (7.9) 30.0 (8.6) 30.0 (7.4) 28.9 (9.0)
Wine 7.0 (4.5) 8.0 (5.3) 7.9 (4.7) 8.8 (5.7) 10.9 (5.1) 7.7 (5.0) 40.1 (10.3)
@@ 7 Discussion
In this article, we focus on functional data that lie on an unknown low-dimensional manifold. We propose a novel proximity measure that takes the label information into account, based on which the FSML has been developed. Coupled with multivariate classifiers, the FSML induces a family of new functional classifiers. We provide the convergence rates of the estimated coordinate map and the asymptotic optimality of our functional classifier when the multivariate classifier is the $k$-NN classifier. We show outstanding classification performance of the FSML methods compared to existing methods under both synthetic and real data examples.
A notable feature of the FSML is that it works particularly efficient when the intrinsic dimension is low, which serves as a motivation for our methodology. In Section 5, we show that the FSML methods outperform several existing methods under two-dimensional manifold models and a Gaussian process model of dimension four. Because discrete observations with large noise certainly inflate the estimated intrinsic dimension, presmoothing is important in our methodology. Note that the intrinsic dimension is equal to the number of PCs when the manifold is flat (i.e., a linear subspace of $L^{2}(\mathcal{T})$), and thus the low-dimensional intrinsic dimension is, to some extent, in contrast to the infinite number of PCs, the core assumption to show asymptotic perfect classification for functional data (Delaigle and Hall,, 2012; Dai et al.,, 2017; Berrendero et al.,, 2018). In our case, perfect classification is possible if the supports of the distributions $X|Y=y$ are disjoint on the manifold. This situation is not trivial as it seems; see Examples 1 and 2 in Section 2.1.

### v2

@@ 5 Simulation Studies
We conduct simulation experiments to evaluate performance of the functional supervised manifold learning (FSML) procedure on finite samples. To construct a functional classifier, we combine the FSML with three classical classifiers: the $k$-NN classifier (FSMLk-NN), the SVM (FSMLSVM), and the linear discriminant analysis (FSMLLDA). For comparison, we consider several existing functional classifiers: the centroid classifier (CC, Delaigle and Hall,, 2012), the functional quadratic discriminant analysis (FQDA, Delaigle and Hall,, 2013; Galeano et al.,, 2015), the nonparametric Bayesian classifier (NB, Dai et al.,, 2017) based on the kernel density estimate, and the functional deep neural network classifier (FDNN, Wang et al.,, 2023). The code to implement the FDNN is available on the authors’ personal website, while the other methods are reproduced using our code.
We generate data from a few functional manifold models as well as two Gaussian process models:
@@
    Torus: $X_{i}(t)=\{2+\cos(\theta_{i})\}\cos(\phi_{i})\sin(2\pi t)+\{2+\cos(\theta_{i})\}\sin(\phi_{i})\cos(2\pi t)+\sin(\theta_{i})\sin(4\pi t)$, where $(\theta_{i},\phi_{i}|Y_{i}=0)$ follows a uniform distribution on $0<\theta<\phi<2\pi$, and $(\theta_{i},\phi_{i}|Y_{i}=1)$ follows a uniform distribution on $0<\phi\leq\theta<2\pi$.
@@
    Gaussian I: $X_{i}(t)=\xi_{i1}\log(t+2)+\xi_{i2}t+\xi_{i3}t^{3}$, where $(\xi_{i1},\xi_{i2},\xi_{i3})^{\top}|Y_{i}=k\overset{iid}{\sim}N(\mu_{k},\Sigma_{k})$ with $\mu_{0}=(-1,2,-3)^{\top},\mu_{1}=(-1/2,5/2,-5/2)^{\top},\Sigma_{0}^{1/2}=\operatorname{diag}(3/5,2/5,1/5)$ and $\Sigma_{1}^{1/2}=\operatorname{diag}(9/10,1/2,3/10)$.
  5. (v)
    Gaussian II: $X_{i}(t)=\mu_{i}(t)+\sum_{j=1}^{50}Z_{ij}\phi_{j}(t)$, where $\mu_{i}(t)=0$ if $Y_{i}=0$ and $\mu_{i}(t)=1$ otherwise, $Z_{ij}|Y_{i}=k\sim N(0,\sigma_{jk}^{2})$ with $\sigma_{j0}=\exp(-j/6)$ and $\sigma_{j1}=\exp(-j/4)$, $\phi_{1}(t)=1,\phi_{2\ell+1}(t)=\sqrt{2}\sin(2\ell\pi t)$, and $\phi_{2\ell}(t)=\sqrt{2}\cos(2\ell\pi t)$.
Under all models above, the label variable $Y_{i}\overset{iid}{\sim}b(1,0.5)$ corresponding to a binary classification setting. The observed data are $(\widetilde{X}_{ij},Y_{i})$, where $\widetilde{X}_{ij}=X_{i}(t_{j})+\epsilon_{ij}$, for $i=1,\ldots,n$ and $j=1,\ldots,J$. Here, $0=t_{1}<\ldots<t_{J}=1$ are equidistant points on $[0,1]$, and $\epsilon_{ij}\overset{iid}{\sim}N(0,\widehat{V}_{X}/20)$, where $\widehat{V}_{X}$ is the sample variance of $X$ integrated on $[0,1]$.
Model (i), modified from model (iii) in Tan et al., (2024), is a time-warping model exhibiting significant phase variation, which is a classical example of a functional manifold (Srivastava et al.,, 2011; Chen and Müller,, 2012). Models (ii) and (iii) are two connected Swiss rolls (see the left panel of Figure 1 for an illustration) and a torus embedded into a functional space, respectively. These geometric objects are two-dimensional manifold examples commonly used in the literature on manifold learning (Tenenbaum et al.,, 2000; Meilă and Zhang,, 2024). Model (iv) is a low-dimensional Gaussian model adopted from Wang et al., 2024b , while model (v) is a high-dimensional one adopted from Dai et al., (2017).
@@
Model $n$ FSML20-NN FSMLSVM FSMLLDA $C$ FQDA NB FDNN
@@
We generate training data following each of models (i) to (v) above with two sampling settings $(n,J)=(200,50)$ and $(500,100)$, while an independent sample of size 500 is generated for testing. The ridged local linear estimator with the plug-in bandwidth as in Section 3.1 is applied to obtain smooth curves before applying any method. We set $k_{\rm{PCA}}=15$ and apply the nested ten-fold CV procedure described in Section 3.4 to select $\xi$ and $h_{\rm{reg}}$. The intrinsic dimension $d$ is set to true values for models (i) to (iv) and estimated for model (v) by the minimal neighborhood method (Facco et al.,, 2017) summarized in Section B of the Appendix. For any method involving FPCA, we choose the minimal number of PCs to explain at least 95% of the variance. We apply all the methods to each of the settings repeatedly for 200 times, and the misclassification errors are summarized in Table 1.
We see from Table 1 that the CC performs the best under model (i), closely followed by the NB and FSML methods. The FDNN performs poorly under model (i). Under models (ii) to (iv), the FSML methods significantly outperform the benchmark methods. In particular, the FSML methods achieve near perfect classification under model (ii), where the CC and FQDA perform poorly. It is worth noting that, although we assume nothing about the underlying data distribution, our methods still perform the best under the low-dimensional Gaussian model (iv), while FQDA and NB indeed perform better than ours under the high-dimensional Gaussian model (v). These observations indicate that the FSML methods are particularly efficient if the intrinsic dimension is low, while the FQDA and NB are more powerful when the intrinsic dimension is high, as their asymptotic perfect classification results require the number of PCs tends to infinity. A similar observation has also been noticed in Wang et al., 2024b . We also note that the differences among the three FSML methods are tiny, indicating that the classification performance of the FSML is insensitive to the chosen specific multivariate classifier. Finally, we recall from Remark 2 that the strategy of registration and classification does not perform well under the models above; see Section E of the Appendix for details.
Figure 3: The mean misclassification rates ($\circ$) using different values of $\xi$ under all simulation models with $n=200$. The vertical intervals denote the mean $\pm$ standard deviation of the misclassification rates.
To see the effect of different values of $\xi$ on the classification performance of our methods, we perform FSML20NN on each of the simulation models with different values of $\xi$ and report the results with $n=200$ in Figure 3. The corresponding results with $n=500$ are reported in Section E of the Appendix. The parameter $h_{\rm{reg}}$ is chosen via the ten-fold CV and other parameters are set the same as before. Figure 3 shows that, for models (ii) and (iii), the mean misclassification rates achieve well-separated minima at $\sqrt{\xi}=20$, while for other models, the mean misclassification rates drop at first but fluctuate little with larger $\xi$. The latter phenomenon is mainly because once the value of $\xi$ is higher than a threshold, the patterns of the manifold learning outcomes stay similar, and thus the classification performance of any classifier does not change much. The reason for presenting $\sqrt{\xi}$ instead of $\xi$ is that the scale of the penalty term in (2.1) is roughly proportional to $\sqrt{\xi}$. Performances of FSMLSVM and FSMLLDA are almost identical to that of FSMLk-NN, and thus they are not reported.
@@
One of the applications of the FSML is data visualization. To illustrate this, we generate two samples from models (ii) and (iii), respectively, and apply the FSML, as well as the FPCA for comparison. We show in Figure 4 the smoothed data by the ridged local linear estimator, the first two PCs by the FPCA, and 2-dimensional manifold learning outcomes by FSML with several $\xi$’s. We see that the FSML properly unfold the Swiss rolls and the outcomes can be linearly separated according to their classes (even with $\xi=0$) under model (ii). Similarly, the FSML outcomes from different classes are also separable with $\xi=4$ under model (iii). In contrast, the first two PCs do not linearly separate two classes under both of the models.
@@ 6 Real Data Applications
The first dataset concerns gene expression levels of a yeast cell based on the $\alpha$ factor experiment (Spellman et al.,, 1998, data available at the supplemental material). Several thousand genes were measured every seven minutes from 0 to 119 minutes, of which 612 genes were found to fluctuate periodically within the cell cycle. We focus on the subset that exhibits significant periodic features: the trajectories belonging to the $G_{1}$ group are coded with $Y=0$ and the trajectories belong to the $G_{2}$ and $M$ groups are coded with $Y=1$. Here, the groups $G_{1},G_{2}$ and $M$ correspond to different phases of the cell cycle. For example, $G_{1}$ is known as the first gap phase, during which the cell grows and prepares for DNA replication. We then obtain a sample of 473 trajectories, among which 223 (resp., 250) trajectories belong to $Y=0$ (resp., $Y=1$). Because the observations are quite noisy and the number of observations per gene is moderate, we use the local linear estimator with a manually chosen large bandwidth to recover smooth functional data.
@@
The raw data for these three datasets are presented in Figure 5. To evaluate the classification performance, we perform each of the functional classifiers based on a 5-fold CV on each of the above datasets. The procedure is repeated 100 times, and the mean and standard deviation of the CV misclassification errors in percentage are reported in Table 2. To implement the FSML methods, the number $k_{\rm{PCA}}$ is set to 10, the intrinsic dimension $d$ is estimated by the minimal neighborhood method (Facco et al.,, 2017), and $\xi$ and $h_{\rm{reg}}$ are chosen following the same criterion as in Section 5. In addition, graphs of the misclassification rates of FSML5-NN with different values of $\xi$ are shown in Section E of the Appendix.
We see from Table 2 that, in the yeast gene dataset, the CC performs the best with a little advantage over the FSML methods, followed by the FQDA and NB, while the FDNN performs less well. Under the DTI dataset, the FSML methods clearly outperform the others. Finally, the FSML methods perform the best under the wine dataset, closely followed by the NB. Overall, the FSML methods demonstrate highly competitive performance compared to other approaches under these three datasets.
Table 2: Mean (standard deviation) CV misclassification errors in percentage under all real datasets with the best result highlighted in boldface.
Dataset FSML5-NN FSMLSVM FSMLLDA $C$ FQDA NB FDNN
Yeast gene 5.8 (0.5) 6.9 (0.7) 5.7 (0.6) 5.5 (0.3) 7.5 (0.6) 8.0 (0.6) 15.6 (8.3)
DTI 23.6 (1.6) 24.1 (1.7) 23.6 (1.7) 26.1 (1.7) 30.5 (2.1) 30.2 (2.4) 30.0 (2.0)
Wine 7.0 (1.1) 7.1 (1.2) 7.1 (1.5) 8.7 (1.1) 10.6 (1.5) 7.6 (1.2) 37.6 (2.8)
@@ 7 Discussion
In this article, we focus on functional data that lie on an unknown low-dimensional manifold. We propose a novel proximity measure that takes the label information into account, based on which the FSML has been developed. Coupled with multivariate classifiers, the FSML induces a new family of functional classifiers. We provide the convergence rates of the estimated coordinate map and the asymptotic optimality of our functional classifier when the multivariate classifier is the $k$-NN classifier. We show highly competitive classification performance of the FSML methods compared to existing methods across both synthetic and real data examples.
A notable feature of the FSML is that it works particularly efficient when the intrinsic dimension is low, which serves as a motivation for our methodology. In Section 5, we show that the FSML methods outperform several existing methods under two-dimensional manifold models and a Gaussian model of dimension three. Because discrete observations with large noise certainly inflate the estimated intrinsic dimension, presmoothing is important in our procedure. Note that the intrinsic dimension is equal to the number of PCs when the manifold is flat (i.e., a linear subspace of $L^{2}(\mathcal{T})$), and thus the low-dimensional intrinsic dimension is, to some extent, in contrast to the infinite number of PCs, the core assumption to show asymptotic perfect classification for functional data (Delaigle and Hall,, 2012; Dai et al.,, 2017; Berrendero et al.,, 2018). In our case, perfect classification is possible if the supports of the distributions $X|Y=y$ are disjoint on the manifold. This situation is not trivial as it seems; see Examples 1 and 2 in Section 2.1.
@@
Acknowledgments
We thank the Editor, the Associate Editor and two referees for their careful reviews and insightful comments that greatly improved our manuscript. Tan’s research was supported by the NSFC (No. 12401363 and 12471263), the Fundamental Research Funds for the Central Universities, and the Key Laboratory of Intelligent Computing and Applications (Ministry of Education). Zang’s research was supported by the Young Faculty Research Grant (No. 110051360025XN077-56) of North China University of Technology.
Disclosure Statement
The authors report there are no competing interests to declare.
Supplementary Materials
Appendix:
The Appendix “App_JCGS-25-043.pdf” contains technical proofs, further implementation details, and additional experimental results.
Code and datasets:
The code and datasets can be used to reproduce the numerical results in Sections 5 and 6. Please refer to the “Readme.pdf” file to see the instruction of using them. The code is also available at https://github.com/ruoxut/FunctionalManifoldLearning.